# Wasserstein Proximal Coordinate Gradient Algorithms

### v1

@@ Abstract
Motivated by approximation Bayesian computation using mean-field variational approximation and the computation of equilibrium in multi-species systems with cross-interaction, this paper investigates the composite geodesically convex optimization problem over multiple distributions. The objective functional under consideration is composed of a convex potential energy on a product of Wasserstein spaces and a sum of convex self-interaction and internal energies associated with each distribution. To efficiently solve this problem, we introduce the Wasserstein Proximal Coordinate Gradient (WPCG) algorithms with parallel, sequential and random update schemes. Under a quadratic growth (QC) condition that is weaker than the usual strong convexity requirement on the objective functional, we show that WPCG converges exponentially fast to the unique global optimum. In the absence of the QG condition, WPCG is still demonstrated to converge to the global optimal solution, albeit at a slower polynomial rate. Numerical results for both motivating examples are consistent with our theoretical findings.
@@ 1 Introduction
The task of minimizing a functional over the space of probability distributions is common in statistics and machine learning, with a wide range of applications in nonparametric statistics [29, 52], Bayesian analysis [45, 53, 22, 30], online learning [23, 5], single cell analysis [31] and spatial economies [8]. Since a probability distribution is an infinite-dimensional object with rich geometric structures, analysis of such an optimization problem requires special treatment and usually leverages optimal transport theory where the convergence is characterized through the Wasserstein metric. In this paper, we consider the problem of jointly minimizing an objective functional over $m$ distributions $\{\rho_{j}\}_{j=1}^{m}$, where $\rho_{j}\in\mathscr{P}_{2}(\mathcal{X}_{j})$, the space of all probability distributions over the $j$-th (Euclidean) domain $\mathcal{X}_{j}$ with finite second-order moments. Precisely, we consider the following general optimization problem:
@@
Example 1.1 (Mean-field inference in variational Bayes).
@@
We use $\Pi_{n}$ to denote the corresponding posterior distribution. In practice, $\pi_{n}$ is often computationally intractable due to the lack of an explicit form of the normalizing constant, i.e., the denominator in (3). Mean-field variational inference (MFVI) [7], which approaches this task by turning the integration problem into an optimization problem, can be formulated as finding a closest fully factorized distribution $\widehat{\pi}_{n}=\bigotimes_{j=1}^{m}\widehat{\rho}_{j}$ to approximate the target posterior distribution $\Pi_{n}$ with respect to the KL divergence; that is, computing
@@
where $U_{n}(\theta)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}\,|\,\theta)$ corresponds to the (averaged) negative log-likelihood function of data $X^{n}$. It is then apparent that above is a special case of the general formulation (1) by taking $V=nU_{n}-\log\pi$, $h_{j}(x)=x\log x$, $W_{j}=0$, and $\mathcal{X}_{j}=\Theta_{j}$. The optimal solution $(\widehat{\rho}_{1},\dots,\widehat{\rho}_{m})$ corresponds to the mean-field approximation of the joint posterior distribution of the parameter $\theta=(\theta_{1},\dots,\theta_{m})$ via the relationship $\widehat{\pi}_{n}=\bigotimes_{j=1}^{m}\widehat{\rho}_{j}$. Ghosh et al. [22] shows the connection between MFVI (4) and the objective functional (1) without proposing practical algorithm for solving (4). Yao and Yang [53] focuses on solving special cases of problem (4) with $m=2$ blocks corresponding to continuous model parameters and discrete latent variables. ∎
Example 1.2 (Equilibrium in multi-species systems with cross-interaction).
Multi-species systems [13, 16] arise in applications in cell biology [39] and population dynamics [55]. In this example, we consider the following non-local multi-species cross-interaction model with diffusion,
@@
In particular, $\rho_{j}(\cdot,t)$ corresponds to the density function associated with the distribution of $X_{j}(t)$ with initialization $X_{j}(0)\sim\rho_{j}(\cdot,0)$ for all $j\in[m]$. Another common class of $\mathcal{H}_{j}$ has $h_{j}(\rho_{j})=\rho_{j}^{m_{j}}$, $m_{j}>1$, corresponding to diffusion in porous media [4, 46]. For such an entropy functional $\mathcal{H}_{j}$, there is no analogous stochastic differential equation (SDE) corresponding to equation (5). For symmetric multi-species models where $K_{ij}=K_{ji}$, finding the equilibrium of equation (5) is equivalent to solving the optimization problem (1) with $V(x)=\sum_{i=1}^{m}V_{i}(x_{i})-\sum_{1\leq i<j\leq m}K_{ij}(x_{i}-x_{j})$ and $W_{j}(x_{j},x_{j}^{\prime})=-K_{jj}(x_{j}-x_{j}^{\prime})/2$ (cf. Proposition 5.3). The symmetric interaction kernel assumption is natural in physics applications due to Newton’s third law of motion.
The demand of computing the stationary distribution (or equilibrium) of a multi-species system naturally arises in physical science, such as chemical engineering [12, 38]. However, most existing literature considers computational methods for calculating the equilibrium of interacting particle systems with only one species. For example, [24] consider the case where $m=1$, $V=h=0$, and the interaction potential $W$ has attractive-repulsive power-law form, by approximating the equilibrium measure with a series of orthogonal polynomials. Under their approximating schemes, the original problem of finding the equilibrium turns into an optimization problem of solving the coefficient of each polynomial. In the multi-species setting, [37] studied the equilibrium of multi-species fractional reaction-diffusion systems by simply approximating the spatial derivatives with second-order Taylor expansion without analyzing the approximation error of their methods. ∎
@@ 1.1 Our contributions
In this paper, we propose the Wasserstein Proximal Coordinate Gradient (WPCG) algorithm as a tool to minimize a composite geodesically convex functional of form (1) over multiple distributions, which extends the coordinate descent algorithm from the Euclidean space to the space of probability distributions. We provide detailed convergence analysis for WPCG with three updating schemes: parallel, sequential and random. Specifically, we show that: (i) under the condition that the potential function $V$ is smooth and convex, and $\{h_{j}\}_{j=1}^{m}$ and $\{W_{j}\}_{j=1}^{m}$ are convex, WPCG converges to a global optimal solution of (1) at rate $O(1/k)$, where $k$ denotes the iteration count; (ii) under the additional quadratic growth (QG) condition in (17), WPCG converges to the unique global optimum exponentially fast in the Wasserstein-2 metric. Note that QG condition is a weaker requirement than the strong convexity on $V$ and/or $\{h_{j}\}_{j=1}^{m}$ and $\{W_{j}\}_{j=1}^{m}$; see Figure 1 for the implications of various assumptions for proving convergence in optimization and Section 3 for further details. In addition, implementation of WPCG with the parallel updating scheme inherently supports parallelization, which enhances its computational scalability to high-dimensional problems.
Previous studies, including [2, 41, 49], predominantly concentrate on using Wasserstein gradient flow for the minimization of a functional over a single distribution. To the best of our knowledge, the current work is among the first study to:
@@
A second challenge in demonstrating the convergence of WPCG lies in the difficulty of proving a Polyak–Łojasiewicz (PL) type inequality, which requires some positive power of certain norm of the “Wasserstein gradient” of target functional $\mathcal{F}$ at any $\rho_{1:m}:=(\rho_{1},\ldots,\rho_{m})$ to be lower bounded by a multiple of $\mathcal{F}(\rho_{1:m})-\mathcal{F}(\rho^{*}_{1:m})$, where $\rho^{*}_{1:m}$ denotes a global minimizer of $\mathcal{F}$. It is known [28] that proving a PL inequality is crucial and common for analyzing gradient-based optimization algorithms in the Euclidean case. A PL inequality is typically implied by strong convexity or a log-Sobolev inequality (LSI). It is also known that a quadratic growth (QG) condition with convexity is strictly weaker than the strong convexity (see discussions after Assumption D). While the QG condition together with convexity implies LSI for the KL-type functional over a single distribution, there is no corresponding result for a general functional over multiple distributions $\rho_{1:m}$. The main technical difficulty comes from the lack of tensorization for a “multivariate” Wasserstein gradient.
@@ 1.2 Related work
Optimization over one distribution. The seminal paper by Jordan, Kinderlehrer, and Otto [27] introduces an iterative scheme (now commonly known as the JKO scheme) serving as the time-discretization of a continuous dynamic in the space of probability distributions following the direction of the steepest descent of a target functional with respect to the Wasserstein-$2$ metric. In later developments, the JKO scheme has been recognized as a promising numerical method for optimizing a functional over a probability distribution and can be viewed as a special proximal gradient update [40] relative to the Wasserstein-$2$ metric in the space of probability distributions (see Section 2.2 for more details). Therefore, in the remainder of the paper, we will primarily use the term Wasserstein proximal gradient scheme instead of the JKO scheme, as this terminology more clearly suggests its close link with general convex optimization.
@@
Recently, several other discretization methods are proposed for minimizing a functional over a single distribution (i.e., $m=1$ in our setting). Different from the Wasserstein proximal gradient scheme that can be viewed as an implicit (backward) scheme for discretizing the Wasserstein gradient flow (WGF), some explicit (forward) schemes, analogous to the usual gradient descent on the Euclidean space, are considered and analyzed for solving specific problems. For example, [15] analyze a gradient descent algorithm for computing the barycenter on the Bures–Wasserstein manifold of centered Gaussian probability measures. One key ingredient in their convergence analysis is to prove a quadratic growth condition for the barycenter functional that leads to a PL inequality and implies the exponential convergence of the algorithm. For the KL divergence functional $\operatorname{KL}(\,\cdot\,\|\,\rho^{\ast})$ with $\rho^{\ast}\,\propto\,e^{-V}$ for some strongly convex and smooth potential function $V$[49] shows that a symmetrized Langevin algorithm, which is an implementable discretization of the Langevin dynamics, reduces the bias in the classical unadjusted Langevin algorithm [19] and attains exponential convergence up to a step size dependent discretization error. [41] consider a more general setting where the objective functional consists of the same potential functional and a general non-smooth term that is convex along generalized geodesics on the Wasserstein space; they propose a forward-backward algorithm, which runs a forward step (gradient descent) for the potential functional and a backward step (proximal gradient) for the non-smooth term, and prove its exponential convergence under the same conditions on potential $V$.
@@
Optimization over special cases of multiple distributions. Several recent works focus on solving an important special case of problem (1) in the context of MFVI as we described in Example 1.1, where a multi-dimensional Bayesian posterior distribution is approximated by the product of several lower-dimensional distributions relative to the KL divergence. Yao and Yang [53] consider a block MFVI for Bayesian latent variable models with two blocks, one for the discrete latent variables and one for the continuous model parameters, which corresponds to problem (1) with $m=2$. They propose and analyze a majorization-minimization algorithm for solving MFVI, which can be viewed as a distributional extension of the classical expectation–maximization (EM) algorithm. Due to the special property that minimizing the latent variable block in the problem admits a closed-form expression, their algorithm is effectively a time-discretized WGF for optimizing a single distribution, with an effective potential function $V$ changing over the iterations. When the population-level log-likelihood function is locally strictly concave, they show that their algorithm converges exponentially fast to the solution of MFVI. Ghosh et al. [22] study a WGF-based algorithm for solving MFVI without latent variables. Their study shows that the discretized flow (Wasserstein proximal gradient scheme) converges to a mathematically well-defined continuous flow when the step size is small, assuming certain conditions on the potential function $V$. However, their analysis is asymptotic and they do not directly examine the convergence of the time-discretized algorithm, leaving it unclear if the discrete dynamic system is stable so that the discretization error does not accumulate over time and whether an explicit convergence rate can be obtained.
@@ 2.1 Wasserstein space and geodesics
[…]
Note that $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is not a flat metric space, but is positively curved in the Alexandrov sense [2]. Geodesics on a curved space are the shortest paths connecting two points. $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is indeed a geodesic space. For any $\mu_{0},\mu_{1}\in\mathscr{P}_{2}^{r}(\mathcal{X})$ and $\gamma\in\Pi_{o}(\mu_{0},\mu_{1})$, the constant-speed geodesic connecting $\mu_{0}$ and $\mu_{1}$ is $\mu_{t}=(\pi_{t})_{\#}\gamma$, where $\pi_{t}=(1-t)\pi^{0}+t\pi^{1}$.

### v2

@@ Abstract
Motivated by approximation Bayesian computation using mean-field variational approximation and the computation of equilibrium in multi-species systems with cross-interaction, this paper investigates the composite geodesically convex optimization problem over multiple distributions. The objective functional under consideration is composed of a convex potential energy on a product of Wasserstein spaces and a sum of convex self-interaction and internal energies associated with each distribution. To efficiently solve this problem, we introduce the Wasserstein Proximal Coordinate Gradient (WPCG) algorithms with parallel, sequential, and random update schemes. Under a quadratic growth (QG) condition that is weaker than the usual strong convexity requirement on the objective functional, we show that WPCG converges exponentially fast to the unique global optimum. In the absence of the QG condition, WPCG is still demonstrated to converge to the global optimal solution, albeit at a slower polynomial rate. Numerical results for both motivating examples are consistent with our theoretical findings.
Keywords: composite convex optimization, coordinate descent, optimal transport, Wasserstein gradient flow, mean-field variational inference, multi-species systems.
@@ 1 Introduction
The task of minimizing a functional over the space of probability distributions is common in statistics and machine learning, with a wide range of applications in nonparametric statistics (Kiefer and Wolfowitz, 1956; Yan et al., 2023), Bayesian analysis (Trillos and Sanz-Alonso, 2020; Yao and Yang, 2022; Ghosh et al., 2022; Lambert et al., 2022), online learning (Guo et al., 2022; Ballu and Berthet, 2022), single cell analysis (Lavenant et al., 2021), and spatial economies (Blanchet et al., 2016). Since a probability distribution is an infinite-dimensional object with rich geometric structures, analysis of such an optimization problem requires special treatment and usually leverages optimal transport theory where the convergence is characterized through the Wasserstein metric. In this paper, we consider the problem of jointly minimizing an objective functional over $m$ distributions $\{\rho_{j}\}_{j=1}^{m}$, where $\rho_{j}\in\mathscr{P}_{2}(\mathcal{X}_{j})$, the space of all probability distributions over the $j$-th (Euclidean) domain $\mathcal{X}_{j}$ with finite second-order moments. Precisely, we consider the following general optimization problem:
@@
Example 1 (Mean-field inference in variational Bayes).
@@
We use $\Pi_{n}$ to denote the corresponding posterior distribution. In practice, $\pi_{n}$ is often computationally intractable due to the lack of an explicit form of the normalizing constant, i.e., the denominator in (3). Mean-field variational inference (MFVI) (Bishop and Nasrabadi, 2006), which approaches this task by turning the integration problem into an optimization problem, can be formulated as finding a closest fully factorized distribution $\widehat{\pi}_{n}=\bigotimes_{j=1}^{m}\widehat{\rho}_{j}$ to approximate the target posterior distribution $\Pi_{n}$ with respect to the KL divergence; that is, computing
@@
where $U_{n}(\theta)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}\,|\,\theta)$ corresponds to the (averaged) negative log-likelihood function of data $X^{n}$. It is then apparent that above is a special case of the general formulation (1) by taking $V=nU_{n}-\log\pi$, $h_{j}(x)=x\log x$, $W_{j}=0$, and $\mathcal{X}_{j}=\Theta_{j}$. The optimal solution $(\widehat{\rho}_{1},\dots,\widehat{\rho}_{m})$ corresponds to the mean-field approximation of the joint posterior distribution of the parameter $\theta=(\theta_{1},\dots,\theta_{m})$ via the relationship $\widehat{\pi}_{n}=\bigotimes_{j=1}^{m}\widehat{\rho}_{j}$. Ghosh et al. (2022) shows the connection between MFVI (4) and the objective functional (1) without proposing practical algorithm for solving (4). Yao and Yang (2022) focuses on solving special cases of problem (4) with $m=2$ blocks corresponding to continuous model parameters and discrete latent variables.
Concurrent and independent to our work, Arnese and Lacker (2024) and Lavenant and Zanella (2024) have also investigated the minimization of (4) with coordinate ascent variational inference (CAVI) using optimal transport theory. Arnese and Lacker (2024) explored the convergence of CAVI with parallel and sequential update schemes, while Lavenant and Zanella (2024) studied the random update scheme. Both papers primarily emphasized the theoretical analysis of the CAVI algorithm, demonstrating similar convergence results to ours for CAVI when the posterior distributions are log-concave. In contrast, our work introduces a novel algorithm that differs from CAVI and studies its algorithmic convergence rate. Furthermore, our analysis is conducted under the quadratic growth condition (detailed in Assumption D), which is weaker than the log-concavity assumption as later shown in Proposition 8. ∎
Example 2 (Equilibrium in multi-species systems with cross-interaction).
Multi-species systems (Carrillo et al., 2018; Daus et al., 2022) arise in applications in cell biology (Pinar, 2021) and population dynamics (Zamponi and Jüngel, 2017). In this example, we consider the following non-local multi-species cross-interaction model with diffusion,
@@
In particular, $\rho_{j}(\cdot,t)$ corresponds to the density function associated with the distribution of $X_{j}(t)$ with initialization $X_{j}(0)\sim\rho_{j}(\cdot,0)$ for all $j\in[m]$. Another common class of $\mathcal{H}_{j}$ has $h_{j}(\rho_{j})=\rho_{j}^{m_{j}}$ with $m_{j}>1$, corresponding to diffusion in porous media (Aronson, 2006; Vázquez, 2007). For such an entropy functional $\mathcal{H}_{j}$, there is no analogous stochastic differential equation (SDE) corresponding to equation (5). For symmetric multi-species models where $K_{ij}=K_{ji}$, finding the equilibrium of equation (5) is equivalent to solving the optimization problem (1) with $V(x)=\sum_{i=1}^{m}V_{i}(x_{i})-\sum_{1\leq i<j\leq m}K_{ij}(x_{i}-x_{j})$ and $W_{j}(x_{j},x_{j}^{\prime})=-K_{jj}(x_{j}-x_{j}^{\prime})/2$ (cf. Proposition 25). The symmetric interaction kernel assumption is natural in physics applications due to Newton’s third law of motion.
The demand of computing the stationary distribution (or equilibrium) of a multi-species system naturally arises in physical science, such as chemical engineering (Carrayrou et al., 2002; Paz-Garcia et al., 2013). However, most existing literature considers computational methods for calculating the equilibrium of interacting particle systems with only one species. For example, Gutleb et al. (2022) consider the case where $m=1$, $V=h=0$, and the interaction potential $W$ has attractive-repulsive power-law form, by approximating the equilibrium measure with a series of orthogonal polynomials. Under their approximating schemes, the original problem of finding the equilibrium turns into an optimization problem of solving the coefficient of each polynomial. In the multi-species setting, Owolabi and Atangana (2019) studied the equilibrium of multi-species fractional reaction-diffusion systems by simply approximating the spatial derivatives with second-order Taylor expansion without analyzing the approximation error of their methods. ∎
@@ 1.1 Our Contributions
In this paper, we propose the Wasserstein Proximal Coordinate Gradient (WPCG) algorithm as a tool to minimize a composite geodesically convex functional of form (1) over multiple distributions, which extends the coordinate descent algorithm from the Euclidean space to the space of probability distributions. We provide detailed convergence analysis for WPCG with three updating schemes: parallel, sequential, and random. Specifically, we show that: (i) under the condition that the potential function $V$ is smooth and convex, and $\{h_{j}\}_{j=1}^{m}$ and $\{W_{j}\}_{j=1}^{m}$ are convex, WPCG converges to a global optimal solution of (1) at rate $O(1/k)$, where $k$ denotes the iteration count; (ii) under the additional quadratic growth (QG) condition in (17), WPCG converges to the unique global optimum exponentially fast in the Wasserstein-2 metric. Note that QG condition is a weaker requirement than the strong convexity on $V$ and/or $\{h_{j}\}_{j=1}^{m}$ and $\{W_{j}\}_{j=1}^{m}$; see Figure 1 for the implications of various assumptions for proving convergence in optimization and Section 3 for further details. In addition, implementation of WPCG with the parallel updating scheme inherently supports parallelization, which enhances its computational scalability to high-dimensional problems.
Previous studies, including (Ambrosio et al., 2005; Salim et al., 2020; Wibisono, 2018), predominantly concentrate on using Wasserstein gradient flow for the minimization of a functional over a single distribution. To the best of our knowledge, the current work is among the first study to:
@@
A second challenge in demonstrating the convergence of WPCG lies in the difficulty of proving a Polyak–Łojasiewicz (PL) type inequality, which requires some positive power of certain norm of the “Wasserstein gradient” of target functional $\mathcal{F}$ at any $\rho_{1:m}:=(\rho_{1},\ldots,\rho_{m})$ to be lower bounded by a multiple of $\mathcal{F}(\rho_{1:m})-\mathcal{F}(\rho^{*}_{1:m})$, where $\rho^{*}_{1:m}$ denotes a global minimizer of $\mathcal{F}$. It is known (Karimi et al., 2016) that proving a PL inequality is crucial and common for analyzing gradient-based optimization algorithms in the Euclidean case. A PL inequality is typically implied by strong convexity or a log-Sobolev inequality (LSI). It is also known that a quadratic growth (QG) condition with convexity is strictly weaker than the strong convexity (see discussions after Assumption D). While the QG condition together with convexity implies LSI for the KL-type functional over a single distribution, there is no corresponding result for a general functional over multiple distributions $\rho_{1:m}$. The main technical difficulty comes from the lack of tensorization for a “multivariate” Wasserstein gradient.
@@ 1.2 Related Work
Optimization over one distribution. The seminal paper by Jordan, Kinderlehrer, and Otto (Jordan et al., 1998) introduces an iterative scheme (now commonly known as the JKO scheme) serving as the time-discretization of a continuous dynamic in the space of probability distributions following the direction of the steepest descent of a target functional with respect to the Wasserstein-$2$ metric. In later developments, the JKO scheme has been recognized as a promising numerical method for optimizing a functional over a probability distribution and can be viewed as a special proximal gradient update (Rockafellar, 1997) relative to the Wasserstein-$2$ metric in the space of probability distributions (see Section 2.2 for more details). Therefore, in the remainder of the paper, we will primarily use the term Wasserstein proximal gradient scheme instead of the JKO scheme, as this terminology more clearly suggests its close link with general convex optimization.
@@
Recently, several other discretization methods are proposed for minimizing a functional over a single distribution (i.e., $m=1$ in our setting). Different from the Wasserstein proximal gradient scheme that can be viewed as an implicit (backward) scheme for discretizing the Wasserstein gradient flow (WGF), some explicit (forward) schemes, analogous to the usual gradient descent on the Euclidean space, are considered and analyzed for solving specific problems. For example, Chewi et al. (2020) analyze a gradient descent algorithm for computing the barycenter on the Bures–Wasserstein manifold of centered Gaussian probability measures. One key ingredient in their convergence analysis is to prove a quadratic growth condition for the barycenter functional that leads to a PL inequality and implies the exponential convergence of the algorithm. For the KL divergence functional $\operatorname{KL}(\,\cdot\,\|\,\rho^{\ast})$ with $\rho^{\ast}\,\propto\,e^{-V}$ for some strongly convex and smooth potential function $V$Wibisono (2018) shows that a symmetrized Langevin algorithm, which is an implementable discretization of the Langevin dynamics, reduces the bias in the classical unadjusted Langevin algorithm (Durmus and Moulines, 2017) and attains exponential convergence up to a step size dependent discretization error. Salim et al. (2020) consider a more general setting where the objective functional consists of the same potential functional and a general non-smooth term that is convex along generalized geodesics on the Wasserstein space; they propose a forward-backward algorithm, which runs a forward step (gradient descent) for the potential functional and a backward step (proximal gradient) for the non-smooth term, and prove its exponential convergence under the same conditions on potential $V$.
@@
Optimization over special cases of multiple distributions. Several recent works focus on solving an important special case of problem (1) in the context of MFVI as we described in Example 1, where a multi-dimensional Bayesian posterior distribution is approximated by the product of several lower-dimensional distributions relative to the KL divergence. Yao and Yang (2022) consider a block MFVI for Bayesian latent variable models with two blocks, one for the discrete latent variables and one for the continuous model parameters, which corresponds to problem (1) with $m=2$. They propose and analyze a majorization-minimization algorithm for solving MFVI, which can be viewed as a distributional extension of the classical expectation–maximization (EM) algorithm. Due to the special property that minimizing the latent variable block in the problem admits a closed-form expression, their algorithm is effectively a time-discretized WGF for optimizing a single distribution, with an effective potential function $V$ changing over the iterations. When the population-level log-likelihood function is locally strictly concave, they show that their algorithm converges exponentially fast to the solution of MFVI. Ghosh et al. (2022) study a WGF-based algorithm for solving MFVI without latent variables. Their study shows that the discretized flow (Wasserstein proximal gradient scheme) converges to a mathematically well-defined continuous flow when the step size is small, assuming certain conditions on the potential function $V$. However, their analysis is asymptotic and they do not directly examine the convergence of the time-discretized algorithm, leaving it unclear if the discrete dynamic system is stable so that the discretization error does not accumulate over time and whether an explicit convergence rate can be obtained.
@@ 2.1 Wasserstein Space and Geodesics
It is known that $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is a metric space called Wasserstein space. Moreover, if $\mathcal{X}$ is complete and separable, $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is complete as well (Bolley, 2008). The infimum in (7) always admits a solution $\gamma$ called an optimal transport plan (Santambrogio, 2015). When $\mu\in\mathscr{P}_{2}^{r}(\mathcal{X})$, the optimal transport plan is unique and takes form $\gamma=(\textrm{Id},T_{\mu}^{\nu})_{\#}\mu$, implying $\nu=(T_{\mu}^{\nu})_{\#}\mu$, where Id is the identity map and $T_{\mu}^{\nu}$ is called the optimal transport map from $\mu$ to $\nu$ (Santambrogio, 2015). Moreover, Brenier’s Theorem (Brenier, 1987) states that this unique optimal transport map $T_{\mu}^{\nu}=\nabla\psi$ is the gradient of a convex function $\psi$.
Note that $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is not a flat metric space, but is positively curved in the Alexandrov sense (Ambrosio et al., 2005). Geodesics on a curved space are the shortest paths connecting two points. $(\mathscr{P}_{2}(\mathcal{X}),\operatorname{W}_{2})$ is indeed a geodesic space. For any $\mu_{0},\mu_{1}\in\mathscr{P}_{2}^{r}(\mathcal{X})$ and $\gamma\in\Pi_{o}(\mu_{0},\mu_{1})$, the constant-speed geodesic connecting $\mu_{0}$ and $\mu_{1}$ is $\mu_{t}=(\pi_{t})_{\#}\gamma$, where $\pi_{t}=(1-t)\pi^{0}+t\pi^{1}$.